OpenStudy (anonymous):

how do you give a power series solution for y''-4y'-5y=0; y(0)=5, y'(0)=0?

OpenStudy (michele_laino):

please try this substitution: \[\large y = c\exp \left( {\lambda x} \right)\] where \[\lambda \] is a parameter, and c is a constant. After that substitution, you should get a quadratic equation for \[\lambda \]

OpenStudy (dan815):

Looks like he wants a power series solution

OpenStudy (michele_laino):

yes! you are right! @dan815 another substitution can be this: \[\large y = \sum\limits_{j = 0}^{ + \infty } {{a_j}{x^j}} \]

OpenStudy (anonymous):

I know that way but the teacher was asking us to solve it using the power series method of y=\[\sum_{n} a_{n}x ^{n}\]

OpenStudy (dan815):